LUCIFER Recalibration after Installation on the Telescope

Executive Summary Block
DAYTIME Tests . . . . . . . . . . . . .
  • LUCI1 re-installation on telescope. done
  • General function tests all but CAM wheel work
  • Internal Focus determination:
  • MOS tests done by Peter, 23Sep2012
  • General reproducibility tests
  • Spectral Alignment Check:
  • FITS header check
  • Field centering and Pupil alignment
  • FlexComp - Calibrate
  • Grating Tilt (Lambda-V) Calibration
  • FlexComp - Verify
  • Offset 0,0 test for MMueller
  • Future ref: Take SpectroSieve flats for calibrations (JKurk)
  • LUCI Aux cryostat - Mask exchange scheduled for 24 Sep

NIGHTTIME Tests . . . . . . . . . . .
  • Twilight Flats (IFF photometric)
  • Orientation on sky
  • Image scale, all cameras:
  • Find Center of Rotation
  • Calibration of AGw transform
  • Field Aberration Measurement
  • Verification of AGw transform
  • PM: Photometric ZPs
  • PM: Field illumination
  • PM: Specphot tests
  • Ghosts, scattered light
  • Run a complete science observation
  • LUCI-to-AGw Focus Measurement
  • Collimation model
  • Pointing model

To recommission LUCI1 after 2012 summer shutdown
Duration and pre-requisites are estimated.
Separated into day-time and night-time tasks

  • WS: Walter Seifert
  • JH: Jochen Heidt
  • PB: Peter Buschkamp
  • DT: Dave Thompson
  • LBTO: Unspecified LBTO personnel


On-telescope day-time (closed dome) calibrations:

LUCI1 re-installation on telescope. Duration 1.0d
Once done, the following are prerequisites before continuing with the recalibrations
  • Wait for LUCI to return to thermal equilibrium - Detector stable @77K
  • Switch on ROE, wait until stable (= time ROE was off, max 2hr)

General function tests Duration: 4h
Responsible: WS
Telescope: @Zenith, Rotator enabled, Lights can be ON
Comments: Use ONLY engineering panels, Main Instrument GUI only AFTER Peter has checked the MOS; start with rotator angle 341deg, repeat for 2-3 other rotator angles

> Basic instrument function tests:
  • Stop SW / Re-start SW (optional, restarted in move to telescope)
  • Repeat at ROT ~ 341, 221, 101 degrees the following:
    • Camera Wheel: Init; move several times between N1.80 & N3.75 camera
    • FW 1 & 2: init, move both wheels front/backwards movements
    • Grating Unit: init, select all gratings/mirror
    • Grating Tilt: set voltages 0V, +-1V
    • Detector focus: move steps to reach neglim & poslim, note # of steps
    • PV: init and move in/out
  • Check that all calib lamps work?

D.Thompson notes...22 Sep 2012 07:30 MST

MCE/ROE software switch was OFF...turned ON <br>
Temps and pressures all look nominal <br>
Full SW shutdown and restart...OK <br>
Message client GUI started to monitor log traffic <br>
TMGUI started, confirmed @ EL=90, ROT=341, TCS alive <br>

@ROT=341: <br>
Cam Eng GUI
   * Initialized, came up at N3.75
   * Move to N1.8 (-747,-10) - N3.75 (747) - N1.8 (-747) - N3.75 (747) - N1.8 (-747) - N3.75 (747)
Filter Wheels: <br>
   * Initialized, came up at blind-blind
   * FW1 moved to P_gam, J_high, H2, OH_1060, Br_gam, FeII, J_low, Y2, clear, Br_gam, FeII, back to blind (+/-194 steps, only one took an extra 10 steps to reach position)
   * FW2 moved to J, Ks, z, clear, zJspec, z,  blind, HKspec, K, H, z, Ks, J, zJspec, clear, back to blind (+/-194 steps, did not see any extra steps taken)
Grating Unit: 
   * Initialized
   * Moved to each grating with no change of voltage
   * 210_zJHK set to -1, 0, +1 volts (untilt, change voltage, tilt, repeat) ... OK
   * 200_H+K set to +1, 0, -1 volts ... OK
   * 150_Ks set to -1, 0, +1 volts ... OK
   * untilt, set back to mirror
Detector Unit:
   * Select N3.75, initialize with negative limit switch, moved -700 then +410 ... OK
   * Select positive limit switch and re-initialize, moved +700 then -220 ... OK
   * Manually find + limit?  Moved +190, 10, 10, 10 positive limit switch now ON (@220 from initialized position)
   * Manually find - limit? Moved -600, -10, -10, -10 negative limit switch now ON (@-410 from initialized position)
   * Select neg. limit switch reference and re-initialize ... OK
Pupil Viewer:
   * Moved to N1.80 camera
   * Initialize PV, move in and out ... OK

Check MOS GUI to make sure masks are locked! <br>
Mask 6 appears to be in the FPU ... will need to move it out for the mask exchange on Monday.

Set rotator to 221, repeat tests above...<br>

@ROT=221: <br>
Cam Eng GUI
   * Started at N1.8, Initialized, came up at 1.8
   * Move to N3.75 (747) - N1.8 (-747,-10) - N3.75 (747) - N1.8 (-747) - N3.75 (747) - N1.8 (-747) 
Filter Wheels: <br>
   * Initialized, came up at blind-blind
   * FW1 moved to P_gam, J_high, H2, OH_1060, Br_gam, FeII, J_low, Y2, clear, Br_gam, FeII, back to blind (+/-194 steps, three took an extra 10 steps to reach position)
   * FW2 moved to J, Ks, z, clear, zJspec, z,  H, K, blind, zJspec, HKspec, z, Ks, J, zJspec, clear, back to blind (+/-194 steps, did not see any extra steps taken)
Grating Unit: 
   * Initialized
   * Moved to each grating with no change of voltage
   * 200_H+K set to +1, 0, -1 volts (untilt, change voltage, tilt, repeat) ... 
      * Did not stabilize at 0V on first two tries (~3mv noise).  
      * Set to +0.2V OK, back to 0V failed.  
      * Set to -0.2V OK, -0.01V OK, 0V OK, -1V OK
   * 210_zJHK set to -1, 0, +1 volts ... OK
   * 150_Ks set to +1, 0, -1 volts ... OK
   * untilt, set back to mirror
Detector Unit:
   * Select N1.8 (where the camera is currently set), initialize with negative limit switch, moved -700, -700 then +370 ... OK
   * Select positive limit switch and re-initialize, moved +700 then -260 ... OK
   * Manually find + limit?  Moved +230, 10, 10, 10 positive limit switch now ON (@220 from initialized position)
      * Timeout exception on first (?) 10 step move...multiple clicks may have been accepted (slow interface)
      * Positive limit now active, re-initialize
      * Moved +230, 10, 10, 10 (+lim not active, position reads 0), 10 (+lim now active, position still reads 0) 
   * Manually find - limit? Moved -600, -10, -10, -10 negative limit switch not active
      * Moved -1, -1, -1, -lim now ON (@-633 from +lim)
   * Select neg. limit switch reference and re-initialize ... OK
Pupil Viewer:
   * Already on N1.80 camera
   * Initialize PV, move in (+lim active at 309 steps) and out (-lim active at -28 steps) ... OK
      * Offset requests were +320 and -350 steps

Set rotator to 101, repeat tests above...<br>

@ROT=101: <br>
Cam Eng GUI
   * Started at N1.8, Initialized, error: can't move in negative direction (already @ -lim)
   * This may be a logic error in the sequence.  It was already in -lim and tried 5 x -15step moves anyway before error msg

#####################################################################
MST 17:23
*** This is my error - I was on the (blue background) PV tab moving motors ***
*** when I thought I was working on the camera.  The following steps were 
*** executed before I realized this.  

   * Manually set +10 steps, initialize, error again (after 1 x -15 step move)
   * Manually set +50 steps, initialize, same error (after 1 x 15 step move)
   * Manually set +700 steps, reached +lim, set counter to 0
   * Manually moved -500 steps, -lim found at -323 steps
   * Set counter to 0, move +322 steps, +lim found at +307 steps.
   * Set counter to 0, move -306 steps, -1 step, -lim found, still reads -306 steps.

   >>> Just realized I was on PV tab...STOPPING IMMEDIATELY!
   >>> Manual motor offsets above were on PV page, "initialize" on camera.

Both the camera and PV show as in their negative limits (PV out, Cam@N1.8)
Cam wheel not initialized (reference switch not active).  Steps between limits 
on PV are the same as at last ROT angle so it seems to be moving OK.  
#####################################################################

Other tests @ ROT=101 not completed:
Filter Wheels: <br>
Grating Unit: 
Detector Unit:
Pupil Viewer:

MST~11:00 
Peter online...after working through some tests from his side, it appears there is some communication problem that needs 
Kai's intervention and he is not immediately available.  We went as far as rebooting the LUCI computer but the problem persists.  
Peter says it is possible that none of the tests done today actually worked, so I will redo them once the communications are re-established.  

Tim checked all connections...looked OK.
Walter did a quick check of the operation and also found no problems.
Re-doing the full functional checkout...

D.Thompson notes...23 Sep 2012 07:30 MST

No other engineering logins seen through NXclient
MCE/ROE software switch is ON <br>
Temps and pressures all look nominal <br>
Restarted only the Telescope Service, TMGUI was not updating. <br>
Message client GUI started to monitor log traffic <br>
Confirmed @ EL=90, ROT=341, TCS alive <br>

@ROT=341: <br>
Cam Eng GUI
   * Looks already initialized, @N1.8 camera, +lim, -lim, and ref all inactive
   * No response when clicking "Initialize Unit" button
   * Move to N3.75...Error! 2012.09.23 13:53:48 s 
error while executing sequence: "camera unit - positive movement from position: 1"
   * Manually move -15 steps (button responds, but no log message)
   * No response when clicking "Initialize Unit" button (repeated 2x)

Full system SW shutdown and restart...

TMGUI updating <br>

@ROT=341: <br>
Cam Eng GUI
   * Found the "start" logging button at bottom, logging now "active"
   * Came up not initialized, as expected from SW restart
   * "Initialize Unit" successful, @N1.8
   * Move to N3.75 (747) - N1.8 (-747,-10,-10,-10,Error)
2012.09.23 14:10:57 s | error while executing sequence: "camera unit - negative movement from position: 2"
   * Move -15 steps manually, GUI tab blinked, no log message, repeat, same response
   * +lim, -lim, and ref all inactive
   * "Initialize Unit" repeated -15 step moves 43 times then failed:
2012.09.23 14:15:14 s | Unable to initialize camera unit!
   * -lim now shows active
   * Move +15 steps manually, -lim now shows inactive, no log message
   * "Initialize Unit" claims to have succeeded, but log indicates error (and no steps):
2012.09.23 14:19:53 s | error while executing sequence: "camera unit - initializing sequence"
   * Move to N3.75 (747) - N1.8 (-747) - N3.75 (747) - N1.8 (-747) - N3.75 (747) - N1.8 (-747) All worked

There are error messages in the SW Control Center bottom window relating to Camera GUI startup.  Could not cut-and-paste.  <br>
After watching more come up, they seem to be related only to drawing the GUIs on screen.

Filter Wheels: <br>
   * Initialized, came up at blind-K
   * FW2 moved to blind
   * FW1 moved to P_gam, J_high, H2, OH_1060, Br_gam, FeII, J_low, Y2, clear, Br_gam, FeII, back to blind (+/-194 steps, only two took an extra 10 steps to reach position, one in each directionn).  All moves completed.
   * FW2 moved to J, Ks, H, clear, zJspec,  blind, HKspec, K, H, z, Ks, J, zJspec, clear, back to blind (+/-194 steps, no moves needed extra steps).  All moves completed. 
Grating Unit: 
   * Initialized
   * Moved to each grating with no change of voltage: mirror - 150_Ks (+252,+252) - 200_H+K (-253) - 210_zJHK (-253,-253)
   * 210_zJHK set to +1, 0, -1 volts (untilt, change voltage, tilt, repeat) ... OK
   * Note: moving to another grating while "tilt"ed gives precondition exception error.  Untilt, move, then tilt.
   * 200_H+K set to -1, 0, +1 volts ... OK
   * 150_Ks set to +1, 0, -1 volts ... OK
   * untilt, set back to mirror
   * Positive moves are 252 steps, negative are 253?
Detector Unit:
   * Select N1.8 (current camera position), initialize with negative limit switch, moved -700 then +370 ... OK
   * Select positive limit switch and re-initialize, moved +700 then -260 ... OK
   * Manually find + limit?  Moved +255, 5 positive limit switch now ON (@260 from initialized position)
   * Manually find - limit? Moved -625, -5, -5 negative limit switch now ON (@-371 from initialized position, -631 from +lim)
   * Select neg. limit switch reference and re-initialize ... OK
Pupil Viewer:
   * Moved to N1.80 camera
   * Initialize PV (-lim, ref active), move in (now +lim, ref) and back out ... OK

Handed over to Peter for MOS checks at ROT=341.

@ROT=221:
Cam Eng GUI
   * Start log update 
   * Came up initialized (@N1.8, PV out)
   * "Initialize Unit" successful, @N1.8
   * Move to N3.75 (747) - N1.8 (-747) - N3.75 (747) - N1.8 (-747) - N3.75 (747) - N1.8 (-747) All worked
Filter Wheels: <br>
   * Start log update 
   * Re-Initialized, FWs at blind-blind
   * FW1 moved to Y1, P_beta, PVlens, OH_1190, HeI, Y1, OH_1060, FeII, J_low, P_gam, clear, Br_gam, FeII, back to blind (+/-194 steps, two took an extra 10 steps to reach position, both + direction).  All moves completed.
   * FW2 moved to zJspec, Ks, K, J, clear, HKspec, H, z, J, zJspec, clear, back to blind (+/-194 steps, no moves needed extra steps).  All moves completed. 
Grating Unit: 
   * Initialized
   * Moved to each grating with no change of voltage: mirror - 210_zJHK - 150_Ks - 200_H+K 
   * 200_H+K set to +1, 0, -1 volts (untilt, change voltage, tilt, repeat) ... OK
   * 150_Ks set to -1, 0, +1 volts ... OK
   * 210_zJHK set to +1, 0 (timed out), 0.20, 0.02, 0 (stabilized, but red), 0 (stabilized, but red), -0.2, -1 volts ... This grating seems to have trouble stabilizing at 0V, noisy (~3ms), other Vs OK. 
   * untilt, set back to mirror
Detector Unit:
   * Select N1.8 (current camera position), initialize with negative limit switch, moved -700, -700 then +370 ... OK
   * Select positive limit switch and re-initialize, moved +700 then -260 ... OK
   * Manually find - limit?  Moved -365, -5, -5, negative limit switch now ON (@-373 from initialized position)
   * Manually find + limit? Moved +625, +5, positive limit switch now ON (@+260 from initialized position, +633 from -lim)
   * Select neg. limit switch reference and re-initialize ... OK
Pupil Viewer:
   * Moved to N1.80 camera
   * Initialize PV (-lim, ref active), move in (now +lim, ref) and back out ... OK
   * Why is the PV 320 (position reads 309) steps to move in and -350 (position reads -310) to move out?

Peter running MOS checks at ROT=221
Peter running MOS checks at ROT=101

@ROT=101:
Cam Eng GUI
   * Log update still running, still @N1.8, PV out
   * "Initialize Unit" failed (3 x -15 step moves.  -lim is active.  
   * Move +15 steps and -lim off.
   * Initialize fails again (3x -15 steps and fails, -lim now active)
   * Move +50 steps and -lim off.
   * Initialize fails again (9x -15 steps and fails, -lim active again)
   * Move +50 steps 
   * Move 4x -15, -lim again active.
   * Peter suggests:
      * +400 from -lim and init...failed (did not detect N1.8 position?)
      * +900 from -lim and init...failed for same reason, -lim active
      * From -lim went +900, +700, 6x +50 and still no +lim
      * -lim...+12...N1.8...+747...N3.75...+637...N30=+lim  (1396 steps)
      * Return to ROT=341
      * -500, -500, 5x -100, 3 x -500, no -lim reached
      * 4x +500, no +lim reached
      * Init...says it worked @N30, no +lim active
      * move manual -100, to get away from N30, init, claims it was successful @N30
      * Move to N3.75, fails?  Init, gives an error but no red messages
      * Log out and restart SW
      * Init comes back immediately on N30.  
      * N3.75 gives error on negative movement from pos. 3
      * Time to get Walter and Michael involved.  

Other tests @ ROT=101 suspended.  
Walter later clarified that N30!=+lim, +lim just past bottom of N30 detent so it "falls back" once +lim is reached

MST 12:45 Walter back, checking things.

See IT#4240

GEIRS/rsync stuck from 20120809...seems /newdata was mounted readonly. Called SHooper, fixed.

GEIRS also giving an error on obtaining WCS keywords from the TCS. TCS is up and running, LUCI's telescope GUI is updating normally. Called TSargent. Seems to be a single keyword TELMODE='ukown'. I think this will go away as soon as we are not in a preset cancelled state on the IIFGUI.

Readout of the detector looks normal.

Walter said to complete the non-camera tests @ ROT=101

Currently @ROT=341, check cam position
Filter Wheels: <br>
   * Start log update 
   * Re-Initialized, FWs at Br_gam-blind 
   * Br_gam + clear, take pic ... N1.8 well centered.  
   * Set back to blind-blind

And now @ ROT=101:
   * Re-Initialized, FWs at blind-blind 
   * FW1 moved to Y1, P_beta, PVlens, H2, Br_gam, Y1, OH_1060, FeII, J_high, P_gam, clear, Br_gam, FeII, back to blind (+/-194 steps, two took an extra 10 steps to reach position, one each direction).  All moves completed.
   * FW2 moved to zJspec, Ks, K, J, zJspec, clear, HKspec, H, z, J, clear, back to blind (+/-194 steps, 1 move needed an extra 10 steps).  All moves completed. 
Grating Unit: 
   * Initialized
   * Moved to each grating with no change of voltage: mirror - 150_Ks - 210_zJHK - 200_H+K 
   * 200_H+K set to +1, 0, -1 volts (untilt, change voltage, tilt, repeat) ... OK
   * 210_zJHK set to -1, 0, +1 volts ... OK 
   * 150_Ks set to +1, 0, -1 volts ... OK
   * untilt, set back to mirror
Detector Unit:
   * Select N1.8 (current camera position), initialize with positive limit switch, moved +700, -260 ... OK
   * Select negative limit switch and re-initialize, moved -700, -700, then -260 ... OK
Pupil Viewer:
   * Moved to N1.80 camera
   * Initialize PV (-lim, ref active), move in (now +lim, ref) and back out ... OK

Functional checkout complete except for CAMERA wheel problem. LUCI Team working on it.
Summary (repeated at ROT ~ 341, 221, 101 degrees):
    • Camera Wheel: communication problem, under investigation
    • FW 1 & 2: no problems -- NOTE: cable swap for camera made FW2 not move to where you think it did.
    • Grating Unit: no problems
    • Grating Tilt: set voltages: 210_zJHK had trouble locking at 0V, everything else OK
    • Detector focus: no problems
    • PV: no problems

Internal Focus determination: Duration: 0h
Responsible: WS
Telescope: @Zenith, Rotator @ 341, Lights can be ON
LUCI: Optic Sieve mask
Done as part of the general function tests, listed here to record results

> Detector stage focus calibration:
  • See ~engineer/TemplateScripts/focusN180.txt and focusN375.txt
  • adjust step range if needed
  • Insert best focus values for both cameras into detector unit GUI config
  • Check tilt of detector w.r.t. focal plane in these data?

Best Focus values:
  • N1.80 = 370 (focus 320 400 10)
  • N3.75 = 370 (focus 330 430 10)

Both focus curves are asymmetric, flatter to shorter step counts. Best focus values are by eye, not from a fit to the data.

To update:
  • Open FOCUS engineering GUI
  • Edit configuration
  • N375
    • N_3_75_REFERENCE_OFFSET_FROM_NEGLIM change from 410 to 370
    • N_3_75_REFERENCE_OFFSET_FROM_POSLIM change from -220 to -260
    • Yes, these values are the same as for the N180 data!
  • N180 no change from current configuration
  • Click "OK" to save changes
  • Initialize and check that the new values are used.

UT20120926 DThompson  luci1.20120926.NNNN

Focus check w/BrG N3.75 best focus @350?  

Configure K, N3.75, focus 330 430 10
   * 0001 Dark 
   * 0002 -  0012 focus sequence - best focus @370 (fit actually 367.5, asymmetric)
Configure K, N1.80, focus 290 390 20 --> 320 400 10
   * 0013 - 0018 focus sequence too wide a spacing
   * 0019 - 0027 focus sequence 

Re-config and init for N3.75 to check new value
   * 0028 @370 focus mean of 16 central spots 0.998 pixels FWHM, IQ across detector looks good.

MOS tests Duration: 4h
Responsible: PB
Telescope: @Zenith, Rotator enabled, Lights can be ON

> Basic MOS checkout:
Tests not detailed. Peter's steps to complete the MOS checkout should be documented for future reference.

 #############################################################################
 # Once the MOS tests have been completed (and passed!) and both Walter and  #
 # Peter have given their OKs to continue, you can use the main instrument   #
 # GUI for all remaining tests.  Be sure to start the telescope service in   #
 # the LUCI SW control GUI manually.                                         #
 #############################################################################

General reproducibility tests Duration: ?h
Responsible: JH, DT
Telescope: @Zenith, Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam, Spectro Sieve mask.
Comments:
  • move ~5x between the corresponding units
  • Repeat at 3-4 rotator angles (90deg separation)
  • Take image after each move
  • Measure reproducibility

Reproducibility tests:
  • Grating Unit
    • Use arclamp (or arclamp + flatlamp to see continuum?)
    • only moves grating unit between measurements
    • All gratings + N3.75 combos, see table below
    • Closed tilt loop on grating (request nominal CWL)
    • Move between mirror <-> grating 5x
    • take image at each position
  • Cameras
    • Use flatfield lamp or thermal background
    • Only moves camera wheel between measurements
    • Grating Unit fixed on mirror position
    • Move between N1.80 <-> N3.75 5x
    • take image at each position
  • Calculate rms positioning for each set of data

RESULTS
Camera wheel reproducibility test results
   RMS in pixels at ROT=71, 161, 251, 341; Nsamp=5 for all configurations

   CamPos    X___071___Y  X___161___Y  X___251___Y  X___341___Y   Datafile
   N3.75                                                              N375_cam.dat
   N1.80                                                              N180_cam.dat

   Camera test repeated at ROT=341 four times, so Nsamp=20
   CamPos    X___341___Y   Datafile
   N3.75                             N375_at341_cam.dat
   N1.80                             N180_at341_cam.dat

Grating wheel reproducibility test results
   RMS in pixels at ROT=71, 161, 251, 341; Nsamp=5 for all configurations

   GratPos   X___071___Y  X___161___Y  X___251___Y  X___341___Y   Datafile
   Mirror                                                              Mirror_grat.dat
   210_zJHK                                                         210_grat.dat
   200_H+K                                                         200_grat.dat
   150_Ks                                                            150_grat.dat

Datafiles attached to this page as: 

Notes from the data collection:
Data taken for these tests @ ROT=71
All @ clear+K, 2.2um CWL, mirror=4s, gratings=30s, Ar lamp on
Mirror          
210_zJHK          
200_H+K          
150_Ks          

Data taken for these tests @ ROT=161
All @ clear+K, 2.2um CWL, mirror=4s, gratings=30s, Ar lamp on
Mirror          
210_zJHK          
200_H+K          
150_Ks          

Data taken for these tests @ ROT=251
All @ clear+K, 2.2um CWL, mirror=4s, gratings=30s, Ar lamp on
Mirror          
210_zJHK          
200_H+K          
150_Ks          

Data taken for these tests @ ROT=341
All @ clear+K, 2.2um CWL, mirror=4s, gratings=30s, Ar lamp on
Mirror          
210_zJHK          
200_H+K          
150_Ks          

Data files:
  • N1.80:
  • N3.75:


Spectral Alignment Check: Duration: 0h
Responsible: DT, JH
Note: uses data from Grating Unit reproducibility test above?

### Would this be better done with a flat lamp instead of an arc lamp?

  • For each grating+camera combo in table below:
    • Make sure grating tilt stabilization is ON
    • Set to grating to nominal CWL, take dispersed image
    • Set to mirror, take image
    • Only 1 pair of exposures needed for each combination
  • Measure spectrum tilt along the rows
  • Measure vertical (y) displacement between image and spectrum

Define:
  • Ytilt = Ymax - Ymin along dispersed spectrum near Y~1024
  • Yoffset = Yimg - Yspec for a centered pinhole @ X~1024

Data:
Spectral Alignment Ytilt Yoffset Img#
N1.80 + 210_zJHK      
N1.80 + 200_H+K      
N1.80 + 150_Ks      

FITS header check Duration: 0h (offline)
Responsible: JH, DT
  • Check the FITS headers for any 'strangeness' and note any other issues

From April 2012 tests: Need to update (for N3.75 camera):
  • PIXSCALE = 0.11782
  • CRPIX1 = 1010.58
  • CRPIX1 = 1047.14
  • SECPIX1 = -0.11782
  • SECPIX2 = 0.11782
  • CDELT1 = 3.272778e-05
  • CDELT2 = 3.272778e-05
  • CD matrix as appropriate
Note apparently no status variable for flexure compensation (ON/OFF)!
  • M4M*POS not an indicator (FC ON but these numbers are not changing)
  • Mirror 4 (M4) not the unit that does flexure compensation?

Field centering and Pupil alignment Duration: 5h
Responsible: WS
Telescope: @Zenith, Rotator enabled
LUCI: Imaging, N3.75, K band

> Field centering procedure using FM4
  • Take an image and save it
  • Measure gaps at edges of image
  • FM4: move the mirror in steps to center, notes:
    • Enlarge CompensationMirrorGUI to full screen, use right side of it.
    • M1 -> y, M2-> x; + x steps -> moves image to the right, +y to the top
    • 100 steps corresponds to approximately to 1pix
  • Take images with both cameras, verify centering

> Pupil alignment procedure:
  • Take image to show present alignment of the pupil
  • Adjust FM1 + FM4 (WS: document procedure for future reference)?
  • Repeat until alignment is acceptable

luci1.20120926.0029 taken with PV in the closed dome.  Pupil looks good as near as I can tell.  Will repeat on sky when possible.

Flexure Compensation - calibration data Duration: 8h
Responsible: DT, JH
Telescope: EL and Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam
Scripts located in ~observer/LBTO/FlexComp_Calib/

> Flexure compensation calibration:
  • Consider 15x36d or 18x30d steps over full -90 to +450 range
  • Exposure times need to be checked (4s for img, 20s for spec)!
  • MANUALLY TURN FLEXURE COMPENSATION OFF to measure!
  • Best sequence:
    • one config
    • all ROT positions (script does this, ROT starts/ends @ -88d)
    • 3 EL's: 90, 55, 20 degrees (order does not matter)
    • next config
  • Use AR lamp for the grating measurements!
  • Process data following ~/LBTO/Flexures/FlexureCheck_Luci1.pdf
  • Select same source in all three datasets for one config!
  • Rename object_0001.tbl like: N180_G210_E90.tbl (see rename.sh)
  • Send renamed tables to Walter

Camera Filter Grating Mask Lamp Script Name Status
N3.75 clear K Mirror Spectro Sieve Thermal FC_N375_mirror.cal  
N1.80 clear K Mirror Spectro Sieve Thermal FC_N180_mirror.cal  

Turn on Ar lamp manually!

Camera Filter Grating Mask Lamp Script Name Status
N1.80 HKspec 200_H+K Spectro Sieve Ar FC_N180_200HK.cal  
N1.80 clear J 210_zJHK Spectro Sieve Ar FC_N180_210zJHK.cal  
N1.80 clear Ks 150_Ks Spectro Sieve Ar FC_N180_150Ks.cal  
             
N3.75 clear Ks 150_Ks Spectro Sieve Ar FC_N375_150Ks.cal  
N3.75 clear J 210_zJHK Spectro Sieve Ar FC_N375_210zJHK.cal  
N3.75 HKspec 200_H+K Spectro Sieve Ar FC_N375_200HK.cal  

Script: FC_N180_mirror.cal (done my MEdwards)
Script: FC_N180_150Ks.cal (done my MEdwards)

UT20120928 DThompson @LBT 
Data: luci1.20120927.NNNN.fits 
Getting "Failed to rebuild database "transmissions" using LUCIFER on log messages every 60s.  
Scripts: /home/observer/LBTO/FlexComp_Calib
Flexure Compensation set to OFF when collecting these data!
Processing: /home/observer/LBTO/Flexures (see cp.sh and rename.sh)

Continue getting occasional "grating tilt did not stabilize in 15s" messages, but watching the engineering 
GUI it looks like it stabilizes immediately after this warning message.  

Script: FC_N180_200HK.cal, dark=0001 
EL  Files              midas.tbl            plot.ps 
90  0002-0033
55  0034-0065
20  0066-0097

Script: FC_N180_210zJHK.cal, dark=0098
EL  Files              midas.tbl            plot.ps 
20  0099-0130
55  0131-0162
90  0163-0194

Script: FC_N375_mirror.cal, dark=0195
EL  Files              midas.tbl            plot.ps 
90  0196-0227
55  0228-0259
20  0260-0292

Script: FC_N375_150Ks.cal
Error: Unable to initialize the compensation mirror.  IMGUI now not responsive to clicks!
One of the FC tables appears to be missing for the 150_Ks grating, is this the cause?
90  
55  
20  

Script: FC_N375_200HK.cal, dark=0292
20  0293-0301 - Ar lamp not on, darn it!
20  0302-0333
55  0334-0365
90  0366-0397

Script: FC_N375_210zJHK.cal , dark=0398
90  0399-0430
55  0431-0435 Aborted...AGw1 work takes precedence...not completed.
20  

Note: be sure to use the EL~20 also at EL=0 to avoid problems at horizon

Grating Tilt (Lambda-V) Calibration Duration: 2d
Responsible: JH, DT
Telescope: @Zenith, Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam, LS0.75_450um slit, grating tilt loop closed.

Spectra: grat*
Throughslit-images: ls*
Done for N1.80-camera only

Test table:
Camera Filter Grating Mask Lamp Voltages to test
N1.80 HKspec 200_H+K LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear z 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear J 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear H 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear K 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear Ks 150_Ks LS0.75_450um Ar/Xe/Ne 3 different
N1.80 zJspec 200_H+K LS0.75_450um Ar  

Camera Filter Grating Mask Lamp Voltages to test
N3.75 zJspec 200_H+K LS0.75_450um Ar  

First results (measured position of line at the detector in x, input was lambda of line):
If all is OK, lines should always be at the same position within a few pixel

Grat mue x-pos mue x-pos mue x-pos
200H+K        
200z+J        
210z        
210J      
210H      
210K      
150Ks        

Verification of lambda-to-voltage measurements for zjSpec Duration: ?h

Flexure Compensation - verification data Duration: 4h

Responsible:
Telescope: EL and Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam (OK for imaging too, same temp as dome)
Scripts located in /home/observer/LBTO/FlexComp_Verify/
Comment: Verification, only a subset needed
Prereq: Flexure calibrations in place, flexure compensation ON!

> Flexure compensation verification:
  • MANUALLY TURN FLEXURE COMPENSATION ON to verify FC!
  • MCS GUI -> ROTATORS -> MOTION CONTROL, set max speed to 1 deg/sec
  • Set EL, send LUCI rotator to -69 degrees to start
  • Best sequence:
    • one config
    • all ROT positions (script does this, ROT starts/ends @ -69d)
    • 2 EL's: 38 and 72 degrees (mid-points of measurements)
    • next config
  • Process data following ~/LBTO/Flexures/FlexureCheck_Luci1.pdf
  • Copy results files into ~/LBTO/Flexures/FlexComp_Verify
  • MCS GUI -> ROTATORS -> MOTION CONTROL, set max speed to 5 deg/sec
  • Send resulting data to Walter

Camera Filter Grating Mask Lamp Script Name
N3.75 clear K Mirror Spectro Sieve Thermal FC_N375_mirror.cal

Turn on Ar lamp manually!

Camera Filter Grating Mask Lamp Script Name
N1.80 HKspec 200_H+K Spectro Sieve Ar FC_N180_200HK.cal
N1.80 clear J 210_zJHK Spectro Sieve Ar FC_N180_210zJHK.cal

Other notes: subtract a dark or a lamp-off before processing. Should see <~2 pix residuals.

Data: 
Scripts: /home/observer/LBTO/FlexComp_Verify
Processing: /home/observer/LBTO/Flexures (see cp.sh and rename.sh)
Results: 

Set Rotator max speed to 1 degree/sec.

Flexure Compensation set to ON when collecting these data!
Script: FC_N375_mirror.cal

EL  Files         midas.tbl            plot.ps 

72  0001-  0011   series of 10 4-sec darks
72  0011 - 0038   N375_MIRROR_E72.tbl  N375_MIRROR_E72.ps, CALunit IN (by accident),  dX~4pxl, dY~5pxl 
38  0039 - 0066   N375_MIRROR_E38.tbl  N375_MIRROR_E38.ps, CALunit OUT, dX~4pxl, dY~5pxl

flush detector, take 20-sec dark with blind/blind luci.20120303.0067.fits
CALunit IN, Ar ON

Next set (0024 - 0051) has FC OFF. I had not touched this button on the IM GUI since starting, however, 
although "Set Flexure Comp" shows ON selected, the box above this read
OFF and was grey, not ON and green like it was for the MIRROR scripts. FC appears to have turned off (in the change of configuration?)
Allowed script to finish, after which I manually selected ON for Set Flexure Comp and now it is ON and green throughout script. 

38  303.0068 - 304.0023 !!!FC OFF!!! N180_G200_E38_fcOFF.tbl  N180_G200_E38_fcOFF.ps, dX>~3pxl, dY~4pxl
UT date changed... luci.20120304.0024 ... 
38  0024 - 0051  N180_G200_E38.tbl N180_G200_E38.ps, dX~1.2pxl, dY~1.5pxl
72  0052 - 0079  N180_G200_E72.tbl N180_G200_E72.ps, dX~2.0pxl, dY~1.5pxl

Immediately ran next script, FC_N180_G210, but found that it turned off FC. Aborted first exposure, cntl-c'd script and then set FC ON. 
Re-ran script and verify FC Status is ON. (Changing configuration does apparently turn off FC and it needs to be turned on again manually or through the script).
72  0080 - 0107  N180_G210_E72.tbl N180_G210_E72.ps, dX~ 3 pxl, dY~1.7 pxl  HUGE JUMP in X (~8 pxls) on image 0092. 
38  0108 - 0135  N180_G210_E38.tbl N180_G210_E38.ps, dX~ 2 pxl, dY~1.1 pxl 

Test that LUCI actually sends an offset 0 0 to the TCS Duration: 0.25h
Test for M.Mueller's planetary atmospheres project...I made a script that just sent 0.0 0.0 offsets and watched both the TCS and LUCI log messages. LUCI does NOT bother to send the 0,0 offset at all, and (not surprisingly) nothing is received on the TCS side. I made a second script that alternated sending +1mas and -1mas offsets in Y only (like dithering along a slit) and the offsets are sent by LUCI and the TCS appears to execute them (saw the update guider stage position command). I also ran a script that did +/- 1 arc sec offsets and the logging looks the same as for 1mas. It is still possible that GCS does not bother applying these updates until they reach a certain threshold, but it sure looked to me like it was applying them.

On-telescope night-time on-sky calibrations:

Note:
  • PM = Photometric

Twilight Flats : Duration: 0h (Twilight)
Evenings/Mornings when photometric:
Pre-requisite: clear skies
Twilight Flatfields

  • Recommend keeping the twilight flatfield counts below ~15000 ADU if at all possible
  • Broadband flats need to start immediately after sunset

Orientation on sky: Duration: 2h
Responsible: DT, JH
Pre-requisite: Orientation done before AGw transform calibration!
Pre-requisite: field with good astrometric coords, good seeing (<1")
Note: Can be done on Persson stds during PM night
Note: Will update further after some MOS acquisitions

> Orientation data
  • Configure for imaging with N3.75, Ks
  • Preset to field (Persson standards, any low-lat field)
  • Take image pair, script: /home/observer/LBTO/SciencePrep/imagepair
  • Repeat for ~5 fields
  • Derive PA offset, update LEFTZEROPOINT
  • Update PCSInstrument.conf and RESTART PCS!
  • Repeat observations for ~5 fields
  • Check that orientation is OK

LEFTZEROPOINT updated to -187.814 (from -187.764). Orientation at PAreq=0.00 is ± degrees (NN measurements).

DThompson notes, UT20120929  -  Files: luci1.20120929.NNNN
Condensation on LUCI window, RH=69% and little wind

5x10.0s Ks N3.75
BS9170  0002 0003  PA=0
BS9178  0004 
BS9178  0005 0006  PA=192.81 
BS9178  0007 0008  PA=180

Closed for a cap cloud.  
Re-opened

BS9183 0009 0010 1x4.0s
BS9183 0011 0012 5x10.0s
BS9188 0013 0014 5x10.0s

Stone_C1  0065 0066 5x10.0s Ks N1.80 camera 
BS9108     0071 0072  PA=0
BS9116     0073 0074  PA=180

0075 = on-sky pupil image

Note: systematic residuals in the rotator calibration will show as a consistent rotator offset in MOS mask acquisitions that can be "mined" from the event log after science blocks and the rotator ZPs updated. The systematic dX and dY offsets from MOS alignments should also be a measure of the difference between the pointing origin and the mechanical mask center. This should be true for MODS as well as LUCI.

Image scale, all cameras: Duration: 0h
Responsible: DT, JH
Uses orientation or photometry data so 0h needed
N3.75 scale expected to change because of lens swap
Pre-requisite: fields with good astrometric coords, good seeing (<1")
Note: Can use the orientation data (scale is a byproduct)
Note: scale measured by WS from sieve mask data: N3.75=0.118"/pix, N1.80=0.249"/pix

  • N3.75 = (April 2012 0.11782 ± 0.00005)
  • N1.80 =

Find Center of Rotation Duration: 3h
Responsible: DT, JH
Pre-requisite: Any dense field
Use pointing mask to also measure mechanical center of masks

> Rotator center
  • Configure for imaging with N3.75, Ks
  • Preset to any dense star field at PA=0
  • Allow to collimate
  • Take image
  • Preset to same field at PA=180
  • Allow to collimate
  • Take image
  • Also do PA=90, 270
  • Repeat on at least two different fields

> Mean coords of matched stars gives rotator center > Mean coords of reference circle gives mechanical center

  • From April 2012
    • ROTCEN @ 1010.58 1047.14
    • MASKCEN @ 1019.87 1032.50
    • ROTCEN -> MASKCEN +1.09"x -1.73"y (DETXY coords)
    • The latter offset will be the typical MOS mask alignment offset needed (this is my prediction).

DThompson notes, UT20120929  -  Files: luci1.20120929.NNNN
Condensation on LUCI window, RH=80% and little wind.  We upped the air feed to 15psi (was about 7psi)

Upped air pressure to 30, RH also dropping.  SW folks running some tests in the meantime.  

Stone_C1 TRACK mode, FC on, 5x4.0s N3.75
N       PA
0021 270
0022 180
0023   90
0024     0

0021-0023 ROTCEN @ 1015.52 1032.87
0022-0024 ROTCEN @ 1015.57 1033.69            ROTCEN(N3.75) @ 1015.55 1033.28

Stone_C1 TRACK mode, FC on, 5x4.0s N1.80
N       PA
0067  0
0068  90
0069  180
0070  270

0067-0069 ROTCEN @ 1012.07 983.31
0068-0070 ROTCEN @ 1010.50 983.66            ROTCEN(N1.80) @ 1011.28(0.81) 983.49(0.19)

Calibration of AGw transform Duration: 6h
Responsible: DLMiller, JStorm (DT, JH to take data?)

See: http://wiki.lbto.org/twiki/bin/view/Commissioning/TransformationData

Data taken UT20120929 running transform_collect.
  • Data in ~LBTO/AGW_Data/20120929/transform
  • Doug reports the residuals are high (160um rms).

Field Aberration Measurement
Responsible: DLMiller (DT, JH to take data?)

See: http://wiki.lbto.org/twiki/bin/view/Commissioning/FieldAberrationMeasurementDetails

UT20120930 03:12 running field_collect.
  • Doug on the phone.
  • Data in ~LBTO/AGW_Data/20120930/field

Verification of AGw transform Duration: 1h
Responsible: ?LBTO (DT, JH to take data)

Once transform is in place, take another set of data to verify. Note: a repeated dice-5 pattern should be well-aligned with rows/columns

Results from centroids of the 10 samples at each position:
From April 2012
Xcen(rms) Ycen(rms)
1023.420(0.431) 1518.292(0.277)
1021.860(0.468) 501.285(0.251)

20120930 luci.20120930.0016 - 0065
  • 0.55 degree rotation wrt detector
  • 0.6 pixels (70mas) rms positioning uncertainty

*20120930 luci.20120930.0138 - *
  • After intermediate (poor quality, 160um rms) fit
  • 0.XX degree rotation wrt detector
  • 0.X pixels (YY mas) rms positioning uncertainty

20120930 luci.20120930.0163 - 0212
  • Slightly better fit with flexure correction from pointing mask, NN microns rms)
  • 0.XX degree rotation wrt detector
  • 0.X pixels (YY mas) rms positioning uncertainty

PM: Total system throughput and photometric ZPs: Duration: 7h
Responsible: DT, JH
In all BB-filters
Pre-requisite: photometric conditions

### Standard with a range of colors observed at a range of airmass to
### derive ZP@airmass=1 and extinction, perhaps also color terms.

> Photometric zero points
  • Preset to series of photometric standards (AIRM 1.0…2.0)
  • Take imagepair through each BB filter (K, Ks, H, J, z)
  • Minimum ~5 samples per filter spread in airmass, 10 is better

ID EL J EL K EL J

From April 2012:
  • J = 25.049 - 0.096(AIRM)
  • K = 24.446 - 0.052(AIRM)

PM: Field illumination: Duration: 3h
Responsible: DT, JH
N3.75 camera due to new field lens
Pre-requisite: photometric conditions
3h should be enough to do a couple of filters
See /home/observer/LBTO/SciencePrep/Illumination for generic scripts

> Field Illumination
  • Configure for imaging in N3.75 camera
  • Select filter, target (any Persson standard, other fields?)
  • Test DIT, configure NDIT to allow collimation
  • Center source on detector
  • Run at least two scripts per filter:
    • Ddice5_100as_illum.scr (gets the extreme corners, +/-100")
    • D25x_illum1.scr (covers field, monitors extinction)
  • Repeat for all filters (eventually) BB and NB.

Ghosts, scattered light: Done? Duration: 2h
Responsible: DT, JH
Not a must but nice add-on.
Pre-requisite: bright star
Bad weather filler

> Ghost check
  • Configure for imaging, J band, N3.75 camera
  • Point to bright star (any in M5 catalog, pref. not K giants)
  • Offset to put the source in the center of the top-left quadrant
  • Take a 7x4.0s exposure
  • Offset DETXY dy=-120" to place in bottom-left quadrant
  • Take a 7x4.0s exposure
  • Difference to check for ghosting, scattered light

PM: Spectrophotometry tests of the new detector and readout modes Duration: 3h
Responsible: DT, JH
In principle not necessary since RON is known, but since spectroscopy is the dominant observing mode it might be wise to check.
Pre-requisite: spectrophot. std

> Specphot tests of new detector
  • Configure for specphot (wide slit)
  • Preset to any spectrophotometric standard
  • Take data over all major modes (gratings and bandpasses)
  • zJspec at high priority as it is new

Specphot test table
Camera Filter Grating Mask Notes
N1.80 clear K 210_zJHK 3_slit #3 10x30" wide slits

Run a complete science program

LUCI-to-AGw Focus Measurement Duration: 2h
Responsible: DT, JH (+WS)
Pre-requisite: STABLE temps, good (<0.7") seeing

Walter, can you outline how you did this?

Note: Can monitor the seeing from the guide cam during measurements to check for any systematic variations. The guide cam will always be in focus once collimation has converged.

> LUCI-to_AGw Focus Measurement

Note that the DIQ on LUCI images on UT20120930 was not particularly good. Systematically elongated in the N3.75 camera while the guide star looked good.
  • Re-check the internal (detector block) focus.
  • LUCI-to-AGw focus may need to be checked because it is a new AGw.

Still to add to the plan?
  • Update/check the collimation model (will happen in course of other tests)
  • Update/check the pointing model (will happen in course of other tests)

  • Model made with MODS on 20120927 does not work well for LUCI.

-- DavidThompson - 04 Sep 2012
Topic revision: r28 - 09 Oct 2012, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback